# The Effect of Personalization in FedProx: A Fine-grained Analysis on Statistical Accuracy and Communication Efficiency

### v3

@@ Abstract
FedProx is a simple yet effective federated learning method that enables model personalization via regularization. Despite remarkable success in practice, a rigorous analysis of how such a regularization provably improves the statistical accuracy of each client’s local model hasn’t been fully established. Setting the regularization strength heuristically presents a risk, as an inappropriate choice may even degrade accuracy. This work fills in the gap by analyzing the effect of regularization on statistical accuracy, thereby providing a theoretical guideline for setting the regularization strength for achieving personalization. We prove that by adaptively choosing the regularization strength under different statistical heterogeneity, FedProx can consistently outperform pure local training and achieve a minimax-optimal statistical rate. In addition, to shed light on resource allocation, we design an algorithm, provably showing that stronger personalization reduces communication complexity without increasing the computation cost overhead. Finally, our theory is validated on both synthetic and real-world datasets and its generalizability is verified in a non-convex setting.
@@ 1 Introduction
Federated Learning (FL) has emerged as an attractive framework for aggregating distributed data, enabling clients to collaboratively train a shared global model while preserving data privacy. In the currently prevalent paradigm (McMahan et al., 2017), FL is formulated as a finite sum minimization problem focusing on a single shared model. Nevertheless, it has been well recognized that one of the key challenges in FL is the statistical heterogeneity of the client datasets. As the participants collect their own local data, it often reflects client-specific characteristics and is not identically distributed. With high statistical heterogeneity, training a single model for all clients by minimizing their average in-sample loss becomes questionable.
To address this challenge, one solution is to relax the common model constraint and solve alternatively the following objective in FedProx (Li et al., 2020a):
$\min_{{{\bm{w}}^{(g)}},\{{{\bm{w}}^{(i)}}\}_{i\in[m]}}\sum_{i\in[m]}p_{i}\left({L}_{i}({{\bm{w}}^{(i)}},S_{i})+\frac{\lambda}{2}\|{{\bm{w}}^{(g)}}-{{\bm{w}}^{(i)}}\|^{2}\right),$ (1)
where $L_{i}({\bm{w}}^{(i)},S_{i})$ is the empirical risk of the $i$-th client, and $S_{i}$ is the local dataset of the $i$-th client. The set $\{p_{i}\}_{i\in[m]}$ is a collection of weights for the clients satisfying $p_{i}\geq 0$ and $\sum_{i}p_{i}=1$. In (1), each client $i$ learns a local personalized model ${\bm{w}}^{(i)}$ by fitting its local data $S_{i}$, while enabling collaborative learning by shrinking all local models towards a common global one ${\bm{w}}^{(g)}$ with intensity controlled by $\lambda$. The smaller $\lambda$ is, the weaker the coupling of the local models the formulation would enforce thus the higher personalization is. Adjusting the strength of regularization controls the extent to which the ${\bm{w}}^{(i)}$’s is adapted to client $i$’s local data, thus the degree of personalization.
The promising empirical performance of FedProx in learning under heterogeneous data (Li et al., 2020a; Donevski et al., 2021; He et al., 2021; An et al., 2023) has led to a sequence of works studying the problem from a pure computation perspective, aiming at developing efficient algorithms under various problem assumptions to reach a (local) solution (Hanzely & Richtárik, 2020; T Dinh et al., 2020; Li et al., 2021; Acar et al., 2021; Yuan & Li, 2022). However, the fundamental question of how to characterize the accuracy of the models produced by (1) is largely unaddressed. As such, for a given set of clients, it is not even clear if introducing personalization will be beneficial, outperforming the baselines given by training a single global model or pure local training. In addition, communication cost in FL is a critical bottleneck in resource-constrained environments. Therefore, understanding when personalization will improve both training efficiency and statistical accuracy is of paramount importance. We address these questions by providing a fine-grained analysis of the effect of personalization in FedProx. Our contribution is summarized below:
$\bullet$ Minimax-optimal Statistical Accuracy. We analyze the statistical accuracy of the objective in (1), taking statistical heterogeneity into account. For each of the models, we establish a state-of-the-art convergence rate $\mathcal{O}(1/(mn))+\mathcal{O}(1/n\wedge R^{2})$, where $R$ characterizes the statistical heterogeneity and $n$ is the sample size on each client. The first term $\mathcal{O}(1/(mn))$ represents the sample complexity of training a single global model under homogeneity, which is attained by (1) when $R$ is close to zero, highlighting the benefit of FL. For the second term, the first component $\mathcal{O}(1/n)$ is the sample complexity of purely local training, and the second component $\mathcal{O}(R^{2})$ demonstrates the flexibility of FedProx under different statistical heterogeneity, implying the statistical accuracy of FedProx never gets worse than purely local training’s. Such a rate matches the existing lower bound for the problem. Therefore, to the best of our knowledge, this is the first work that demonstrates that with a theoretically-informed adaptive choice of personalization degree characterized by $\lambda$, the objective in (1) achieves a minimax-optimal statistical accuracy across different levels of statistical heterogeneity. The theoretical analysis we use for objective (1) would extend to different objectives in regularization-based personalized federated learning (Smith et al., 2017; Li et al., 2020a; Hanzely & Richtárik, 2020; Hanzely et al., 2020; 2023).
$\bullet$ Provably Enhanced Communication Efficiency. As the objective (1) could be reformulated as a bi-level problem, we propose an algorithm for (1) with an explicit computation cost assessment. We establish the communication complexity as $\mathcal{O}(\kappa\frac{\lambda+\mu}{\lambda+L}\log\frac{1}{\varepsilon})$ and the local computation complexity in each communication round as $\mathcal{\widetilde{O}}({\frac{\lambda+L}{\lambda+\mu}})$. With higher personalization degree, communication complexity would decrease and the computation complexity per round would increase, demonstrating that there is a trade-off between computation and communication costs. Our analysis concludes that model personalization can reduce communication overhead without incurring additional computation expenses.
@@
$\bullet$ Empirical Validation. We conduct numerical studies on logistic regression with synthetic data and real-world data under our problem assumptions. The results corroborate our theoretical findings. We further test the numerical performance of (1) for training CNNs, showing empirically the results, although proved under the convexity assumption, generalize to nonconvex losses.
@@ 2 Related Works
In recent years, federated learning has become an attractive solution for training models locally on distributed clients, rather than transferring data to centralized data centers (Mammen, 2021; Wen et al., 2023; Beltrán et al., 2023). As each client generates its local data, statistical heterogeneity is common with data being non-identically distributed between clients (Li et al., 2020a; Ye et al., 2023). Given the variability of data in network, model personalization is a natural strategy used to improve learning accuracy for each client. Formulations enabling model personalization have been studied independently from multiple fields. For example, meta-learning (Chen et al., 2018; Jiang et al., 2019; Khodak et al., 2019; Fallah et al., 2020) assumes all models follow a common distribution. By minimizing the average loss, these methods aim to learn a meta-model that generalizes well to unseen new tasks. Multi-task learning methods (Agarwal et al., 2020; Liang et al., 2020) and transfer learning methods (Li et al., 2022; He et al., 2024a; b) study formulation similar to (1), but the statistical rate is established for either the average of all models or only the target model. Therefore, it is not clear if all clients can benefit from the procedure and learn a more accurate model than purely local training. Representation learning (Zhou et al., 2020; Wang et al., 2024), on the other hand, focuses on a different model structure where the local models can be represented as the composition of two parts, with one common to all clients and the other specific to each client.
@@
In the FL community, personalized FL methods can be broadly divided into two main strands. Different from the works mentioned previously, studies here focus primarily on the properties of the iterates computed by the algorithms. One line of works (Arivazhagan et al., 2019; Liang et al., 2020; Singhal et al., 2021; Collins et al., 2021) is based on the representation learning formulation, where both the algorithmic convergence and statistical accuracy of the iterates have been investigated. Another line of works (Smith et al., 2017; Li et al., 2020a; Hanzely & Richtárik, 2020; Hanzely et al., 2020; 2023), including FedProx, achieves personalization by relaxing the requirement of learning a common global model through regularization techniques. In particular, algorithms and complexity lower bounds specific to the FedProx objective were studied in works (Li et al., 2020a; Hanzely et al., 2020; T Dinh et al., 2020; Hanzely & Richtárik, 2020; Li et al., 2021), see Table 2 for a detailed comparison. These works study the problem from a pure optimization perspective and cannot show personalization improves learning accuracy.
@@
There are only a few recent works we are aware of studying the statistical accuracy of regularization-based PFL in the form of (1). Specifically, Cheng et al. (2023) investigates the asymptotic behavior of the FedProx estimator under an over-parameterized linear regression model; neither a finite-sample rate nor algorithm is provided. Chen et al. (2023b) studies the FedProx estimator under rather strong assumptions like bounded parameter space. By defining a new notion of algorithmic stability, they establish the non-asymptotic sample complexity for FedProx and induce minimax bound. However, even with some overly strong assumptions, their analysis cannot match the statistical lower bound in most cases.
Notation. $[n]$ denotes the set of integers $\{1,2,\cdots,n\}$ for any positive integer $n$. Let $\|\cdot{}\|_{2}$ denote the $\ell_{2}$ vector norm and the Frobenius matrix norm. For two non-negative sequences $a_{n}$, $b_{n}$, we denote $a_{n}=\mathcal{O}(b_{n})$ as $a_{n}/b_{n}\xrightarrow{}{1}$. $\tilde{\mathcal{O}}(\cdot)$ hides logarithmic factors.
@@ 3 Preliminaries
Considering there are $m$ clients, each client has a collection of $i.i.d$ data $S_{i}=\{z_{ij}\}_{j\in[n_{i}]}$ over possibly different data distributions $\mathcal{D}_{i}$. For each client, the ultimate goal is to find the solution that minimizes the local risk, i.e. ${{\bm{w}}_{\star}^{(i)}}\in\arg\min_{{\bm{w}}}\>\mathbb{E}_{z\sim\mathcal{D}_{i}}\ell({\bm{w}},z)$. Here, $\ell({\bm{w}},z)$ denotes the loss function of interest and ${{\bm{w}}_{\star}^{(i)}}\in\mathbb{R}^{d}$ is called the ground truth local model for client $i$. In this paper, we assume the model dimension $d$ is finite. Under such a setup, two commonly adopted approaches are pure local training (LocalTrain) and pure global training (GlobalTrain):
@@
where $L_{i}({\bm{w}},S_{i}):=\sum_{j=1}^{n_{i}}\ell({\bm{w}},z_{ij})/n_{i}$ denotes the empirical risk of client $i$. In LocalTrain, each client individually minimizes its loss, whereas in GlobalTrain, all clients collaboratively minimize their average loss. Compared to LocalTrain, GlobalTrain integrates information from all clients to improve the learning performance of each client. When there are sufficient similarities between the data distribution $\mathcal{D}_{i}$ of different clients, GlobalTrain would be a better choice as more samples are included. However, if $\mathcal{D}_{i}$’s are highly unrelated, LocalTrain will be a more robust choice.
In practice, the data distribution is neither totally the same nor completely unrelated across clients. To handle the wide middle ground of the two extremes, FedProx solves the following problem
$\min_{{{\bm{w}}^{(g)}},\{{{\bm{w}}^{(i)}}\}_{i\in[m]}}\sum_{i=1}^{m}p_{i}\left({L}_{i}({{\bm{w}}^{(i)}},S_{i})+\frac{\lambda}{2}\|{{\bm{w}}^{(g)}}-{{\bm{w}}^{(i)}}\|^{2}\right).$
The key feature of Fedprox is the incorporation of a proximal (regularization) term, which controls the personalization degree. As $\lambda\rightarrow 0$, the FedProx objective increasingly behaves like LocalTrain with higher personalization degree and reduces to $m$ separate LocalTrain problems in the end. On the other hand, as $\lambda\rightarrow\infty$, the local models ${\bm{w}}^{(i)}$ are shrunk towards the global model ${\bm{w}}^{(g)}$ with lower personalization degree and eventually degenerate to optimize a single shared model. Thus, by adjusting $\lambda$, FedProx interpolates between LocalTrain and GlobalTrain, allowing for a flexible choice of personalization degree to achieve higher modeling accuracy. It’s noticeable that this choice of $\lambda$ also affects the optimization process of FedProx. With larger $\lambda$, more frequent communication between clients is expected to ensure faster convergence, as local updates are more tightly coupled to the global model. In contrast, smaller $\lambda$ value allows for more independent updates, reducing the need for communication. In the next few sections, we provide a theoretical guarantee for the intuition provided above.
In Section 4, we analyze when the personalization level, characterized by $\lambda$, would improve the statistical accuracy of both the global model ${\bm{w}}^{(g)}$ and the local models $\{{{\bm{w}}^{(i)}}\}_{i\in[m]}$. This is in sharp contrast to most of the existing works where the error is established for the sum of them. In Section 5, we explore how such a degree of personalization will affect communication complexity.
4 When will Personalization Improve Statistical Accuracy
In this section, we analyze how the choice of $\lambda$, or the degree of personalization, would improve the statistical accuracy of the solution obtained from the FedProx problem (1). Following the existing literature (Li et al., 2023; Chen et al., 2023a; b; Duan & Wang, 2023), we define the statistical heterogeneity as the Euclidean distance between the true local models ${{\bm{w}}_{\star}^{(i)}}$ and the true global model ${{\bm{w}}_{\star}^{(g)}}:=\sum_{i=1}^{m}p_{i}{{\bm{w}}_{\star}^{(i)}}$. Specifically, we adopt the following assumption:
Assumption 1 (Statistical Heterogeneity).
There exists a nonnegative constant $R$ such that we have $\|{{\bm{w}}_{\star}^{(i)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}\leq R^{2},\forall i\in[m]$.
Intuitively, a larger $R$ indicates a greater divergence between the clients’ data distributions, thus stronger statistical heterogeneity, and vice versa. Therefore, in the FedProx objective, the choice of $\lambda$ should be adjusted according to $R$. More discussion about Assumption 1 can be found in Appendix A.1. Next, we introduce three standard assumptions.

### v4

# Understanding the Statistical Accuracy-Communication Trade-off in Personalized Federated Learning with Minimax Guarantees

@@ Abstract
Personalized federated learning (PFL) offers a flexible framework for aggregating information across distributed clients with heterogeneous data. This work considers a personalized federated learning setting that simultaneously learns global and local models. While purely local training has no communication cost, collaborative learning among the clients can leverage shared knowledge to improve statistical accuracy, presenting an accuracy-communication trade-off in personalized federated learning. However, the theoretical analysis of how personalization quantitatively influences sample and algorithmic efficiency and their inherent trade-off is largely unexplored. This paper makes a contribution towards filling this gap, by providing a quantitative characterization of the personalization degree on the tradeoff. The results further offers theoretical insights for choosing the personalization degree. As a side contribution, we establish the minimax optimality in terms of statistical accuracy for a widely studied PFL formulation. The theoretical result is validated on both synthetic and real-world datasets and its generalizability is verified in a non-convex setting.
Personalized Federated Learning, Statistical Complexity, Communication Complexity
@@ 1 Introduction
Federated Learning (FL) (McMahan et al., 2017) has emerged as a promising learning framework for aggregating information from distributed data, allowing clients to collaboratively train a shared global model in a communication-efficient manner. However, a critical challenge arising in the presence of various data across clients is data heterogeneity. In such cases, a single global model may fail to generalize well to all clients, motivating the need for model personalization. PFL enhances FL by learning personalized models tailored to individual clients, and has demonstrated strong empirical performance in various applications, such as driver monitoring (Yuan et al., 2023), and mobile computing (Zhang et al., 2024).
An important question in PFL is determining the degree of personalization, which controls the transition between fully collaborative training and pure local training. A higher degree of collaboration (less personalization) typically requires more frequent information exchange, potentially improving learning accuracy when client data distributions are similar. Conversely, increasing personalization reduces communication costs by prioritizing localized training, but may lead to higher generalization errors due to the limited size of client datasets (Paragliola, 2022). Understanding this trade-off is essential for optimizing model performance under communication constraints.
Most existing works focus purely on the algorithmic perspective of PFL (Lin et al., 2022; Li et al., 2024b; Wang et al., 2022, 2024b). However, the statistical accuracy of the solutions obtained in PFL remains largely unexplored. As a result, the connections between statistical accuracy, communication efficiency, and their trade-offs are not well understood, leaving a theoretical gap in understanding how to select the optimal personalization degree. In this paper, we fill in this gap by providing a fine-grained theoretical analysis of a widely adopted PFL problem formulation given by  (2) that trains simultaneously local and global models. We quantitatively analyze the influence of personalization on both statistical and optimization convergence rates for each of the local models. Specifically, our contribution can be summarized as follows:
$\bullet$ Statistical Accuracy with Optimal Guarantee. We provide a non-asymptotic statistical convergence rate of the solution of Problem (2), revealing how personalization degree influences the statistical accuracy of each local model. In particular, as the personalization degree increases, the rate approaches to that of pure local training, $\mathcal{O}(1/n)$, where $n$ is the sample size per client. Conversely, decreasing the personalization degree allows the models to utilize information across all clients, achieving a rate closer to $\mathcal{O}(1/(mn)+R^{2})$, where $mn$ is the total sample size of all clients and $R$ quantifies the statistical heterogeneity of the local datasets. We further establish the minimax optimality of the derived statistical rates, demonstrating the tightness of our analysis. To the best of our knowledge, this is the first work to achieve such an optimality.
$\bullet$ Communication Efficiency under Personalization. On the optimization aspect, we treat Problem (2) as a bi-level problem and propose (stochastic) algorithms with an explicit characterization of computation and communication complexity. We establish that the communication cost of finding an $\varepsilon$-solution is $\mathcal{O}(\kappa\frac{\lambda+\mu}{\lambda+L}\log\frac{1}{\varepsilon})$ and the computation cost in terms of gradient evaluations is $\mathcal{O}(\kappa\log\frac{1}{\varepsilon})$, where $\lambda$ is a parameter determining the personalization degree. This demonstrates that a smaller $\lambda$, corresponding to a higher degree of personalization, reduces communication complexity without incurring additional computational overhead.
@@
$\bullet$ Accuracy-Communication Tradeoff with Empirical Validation. Building on our theoretical results for statistical and optimization convergence, we quantitatively characterize the trade-off between statistical accuracy and communication efficiency in PFL and discuss practical insights for selecting the personalization degree. We then conduct numerical studies on logistic regression with synthetic data and real-world data under our assumptions. The results corroborate our theoretical findings. We further test the numerical performance with CNN models, showing that the theoretical results, though established under convexity assumptions, generalize to non-convex settings.
@@ 2 Related Works
In recent years, FL has become an attractive solution for training models locally on distributed clients, rather than transferring data to a single node for centralized processing (Mammen, 2021; Wen et al., 2023; Beltrán et al., 2023). As each client generates its local data, statistical heterogeneity naturally arises with data being non-identically distributed between clients (Li et al., 2020; Ye et al., 2023). Given the variability of data in a network, model personalization is an appealing strategy used to improve statistical accuracy for each client. Formulations enabling model personalization have been studied independently from multiple fields. For example, meta-learning (Chen et al., 2018; Jiang et al., 2019; Khodak et al., 2019; Fallah et al., 2020) assumes all local models follow a common distribution. By minimizing the average validation loss, these methods aim to learn a meta-model that generalizes well to unseen new tasks. Representation learning (Zhou et al., 2020; Wang et al., 2024a) focuses on a setting where the local models can be represented as the composition of two parts, with one common to all clients and the other specific to each client. Our work, however, differs in that we focus on a personalized FL setting with a mixture of global and local models. Closely related are multi-task learning methods (Liang et al., 2020) and transfer learning methods (Li et al., 2022; He et al., 2024a, b), but the statistical rate is established for either the average of all models or only the target model. Therefore, it remains unclear how personalization influences the statistical accuracy of each individual local model in these settings.
@@
In the FL community, personalized FL methods can be broadly divided into two main strands. Different from the works mentioned previously, studies here primarily focus on the properties of the iterates generated by the algorithms. One line of work (Arivazhagan et al., 2019; Liang et al., 2020; Singhal et al., 2021; Collins et al., 2021) is based on the representation learning formulation. Another line of work (Smith et al., 2017; Li et al., 2020; Hanzely & Richtárik, 2020; Hanzely et al., 2020; Li & Richtárik, 2024) achieves personalization by relaxing the requirement of learning a common global model through regularization techniques. In particular, algorithms and complexity lower bounds specific to Problem (2) were studied in (Li et al., 2020; Hanzely et al., 2020; T Dinh et al., 2020; Hanzely & Richtárik, 2020; Li et al., 2021), see Table 1 for a detailed comparison. These works study Problem (2) from a pure optimization perspective and have not provided how personalization influences statistical accuracy and, consequently, the trade-off between statistical accuracy and communication efficiency.
@@
There are only a few recent works we are aware of that study the statistical accuracy of regularization-based PFL. Specifically, Cheng et al. (2023) investigates the asymptotic behavior of the (personalized) federated learning under an over-parameterized linear regression model; neither a finite-sample rate nor algorithms are provided. Chen et al. (2023c) studies Problem (2) and establishes a non-asymptotic sample complexity. However, even with some overly strong assumptions, their analysis cannot match the statistical lower bound in most cases. Furthermore, neither of these works reveals the trade-off. While (Bietti et al., 2022) studies the influence of personalization in PFL, their focus is on the trade-off between privacy and optimization error, leaving the accuracy-communication trade-off unexplored.
Notation. Denote $[n]:=\{1,2,\cdots,n\}$. $\|\cdot{}\|_{2}$ denotes the $\ell_{2}$ norm for vectors and the Frobenius norm for matrices. For two non-negative sequences $\{a_{n}\},\{b_{n}\}$, we denote $a_{n}\lesssim b_{n}$ if $a_{n}\leq Cb_{n}$ for some constant $C>0$ when $n$ is sufficiently large. We also use $a_{n}=\mathcal{O}\left(b_{n}\right)$, whose meaning is the same as $a_{n}\lesssim b_{n}$. $\tilde{\mathcal{O}}(\cdot)$ hides logarithmic factors. For two real numbers $a$, $b$, we let $a\wedge b=\min\{a,b\}$ and $a\vee b=\max\{a,b\}$.
@@ 3 Preliminaries
We consider an FL setting with $m$ clients. Each client $i$ has a collection of data $S_{i}=\{z_{ij}\}_{j\in[n_{i}]}$, where $n_{i}$ is the sample size of client $i$, with elements i.i.d. drawn from distribution $\mathcal{D}_{i}$. The data distributions among the clients are possibly heterogeneous. For each client $i$, the ultimate goal is to find a model that minimizes its local risk at the population level, defined as:
${{\bm{w}}_{\star}^{(i)}}\in\arg\min_{{\bm{w}}}\>\mathbb{E}_{z\sim\mathcal{D}_{i}}\ell({\bm{w}},z),$ (1)
where $\ell({\bm{w}},z)$ is a loss function measuring the fitness of ${\bm{w}}$ to data point $z$. We call ${{\bm{w}}_{\star}^{(i)}}\in\mathbb{R}^{d}$ the ground truth model for client $i$. To understand the influence of model personalization, we study the following widely adopted PFL problem (Hanzely & Richtárik, 2020; Mishchenko et al., 2023; Li et al., 2020):
$\min_{\begin{subarray}{c}{{\bm{w}}^{(g)}}\\ \{{{\bm{w}}^{(i)}}\}_{i=1}^{m}\end{subarray}}\sum_{i\in[m]}p_{i}\left({L}_{i}({{\bm{w}}^{(i)}},S_{i})+\frac{\lambda}{2}\|{{\bm{w}}^{(g)}}-{{\bm{w}}^{(i)}}\|^{2}\right),$ (2)
where ${{\bm{w}}^{(g)}}$ and ${{\bm{w}}^{(i)}}$ are respectively the global and $i$-th local model to be learned, $L_{i}({\bm{w}},S_{i}):=\sum_{j=1}^{n_{i}}\ell({\bm{w}},z_{ij})/n_{i}$ is the empirical risk of client $i$ on its local data $S_{i}$, and the last term is a regularization term with parameter $\lambda$ controlling the personalization degree. The set $\{p_{i}\}_{i\in[m]}$ is a collection of nonnegative weights with $\sum_{i=1}^{m}p_{i}=1$.
In Problem (2), each client $i$ learns a personalized model ${{\bm{w}}^{(i)}}$ by fitting to its local data $S_{i}$, while collaboration is achieved by shrinking the local models ${{\bm{w}}^{(i)}}$ towards a common global model ${{\bm{w}}^{(g)}}$. As $\lambda\to 0$, the influence of the global model diminishes, and Problem (2) increasingly behaves like the LocalTrain problem:
@@
where each client independently trains its model using only local data. In this case, Problem (2) is fully decoupled into $m$ separate problems, achieving the maximum degree of personalization. On the other hand, as $\lambda\to\infty$, the regularization term shrinks all local models ${{\bm{w}}^{(i)}}$ towards the global model ${{\bm{w}}^{(g)}}$. This corresponds to reducing the personalization degree and eventually pushes Problem (2) to another extreme given by:
[…]
where the knowledge from all clients is pooled to train a single global model. Adjusting $\lambda$, therefore, controls the degree of personalization among clients and thus affects the statistical accuracy of the resulting solution. Noticeably, the choice of $\lambda$ also potentially affects solving Problem (2) algorithmically in the FL setting. Intuitively, for large $\lambda$, we anticipate more frequent communications among the clients to facilitate collaboration as the local models are more tightly coupled to the global model. In contrast, a smaller $\lambda$ encourages more independent updates and reduces the need for communication, with LocalTrain (cf.(3)) being an edge case where all clients train independently without any communication.
We quantify both the statistical and optimization error of a PFL algorithm applied to Problem (2) by evaluating the Euclidean distance between the algorithm’s output after $t$ communication rounds, ${\bm{w}}_{t}^{(i)}$, and the ground truth local model ${{\bm{w}}_{\star}^{(i)}}$ for each client $i\in[m]$. Let $\widetilde{\bm{w}}^{(i)}$ denote the solution of $i$-th local model in Problem (2). We have the following decomposition:
$\|{\bm{w}}_{t}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\|^{2}\leq 2\underbrace{\|{\bm{w}}_{t}^{(i)}-\widetilde{\bm{w}}^{(i)}\|^{2}}_{\text{optimization error}}+2\underbrace{\|\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\|^{2}}_{\text{statistical error}}.$
In the rest of the paper, we first establish the statistical and optimization convergence rates independently. We then integrate these results to formally provide the trade-off between statistical accuracy and communication efficiency under different levels of personalization.
4 Convergence Rate Analysis
4.1 Effect of Personalization on Statistical Accuracy
In this section, we analyze how the choice of $\lambda$ influences the statistical accuracy of the solution to Problem (2). Recall the ground truth local model ${{\bm{w}}_{\star}^{(i)}}\in\mathbb{R}^{d}$ defined in (1). Next, we introduce the measure of statistical heterogeneity and define the parameter space for the statistical estimation problem. Specifically, we consider estimating ${{\bm{w}}_{\star}^{(i)}}$ from the following parameter space:
$\mathcal{P}(R)=\left\{\{{{\bm{w}}_{\star}^{(i)}}\}_{i=1}^{m}:\left\|{{\bm{w}}_{\star}^{(i)}}-\sum_{i=1}^{m}p_{i}{{\bm{w}}_{\star}^{(i)}}\right\|^{2}\leq R^{2},\forall i\in[m]\right\}.$ (5)
In (5), the statistical heterogeneity is measured as the Euclidean distance between the ground truth local models and their weighted average. A larger $R$ indicates a larger difference among the clients’ local models, hence stronger statistical heterogeneity. Such a measure is commonly imposed in the existing literature (Li et al., 2023; Chen et al., 2023b, c; Duan & Wang, 2023). More discussion about the parameter space $\mathcal{P}(R)$ can be found in Appendix A.1. We assume the model dimension $d$ is finite.
Under such a parameter space, we aim to investigate the statistical accuracy of the solution to Problem (2) measured by $\mathbb{E}\|\widetilde{\bm{w}}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\|^{2}$, where the expectation is taken with respect to the joint distribution.
Remark 1.
Note that some existing literature adopts alternative metrics, such as individual excess risk (Chen et al., 2023c) to measure the statistical error. These two metrics are equivalent under the strong convexity and smoothness conditions on the loss functions.
We study Problem (2) under the following regularity conditions.